# arXiv:2603.00663v1[cs.RO] 28 Feb 2026

### v1

@@ 7 Conclusion
In this paper, we introduced BPRC, a new algorithm to find the optimum for the MT-VRP. We demonstrated that it finds the optimum with less computation time than a baseline, particularly when the agents have small capacities.
@@ Appendix D Generating Additional Feasible Solutions
This section describes the function used to generate the initial incumbent and generate additional feasible solutions when the RMP is infeasible in Section 5. Given a set of disallowed edges $\mathcal{B}\subseteq\mathcal{E}_{\text{tw}}$, GenerateFeasibleSolution attempts to generate a set of tours solving the MT-VRP while traversing no edge in $\mathcal{B}$. We do so using the DFS in Alg. 2, which attempts to quickly find tours without regard to cost. Alg. 2 takes as input a set of disallowed edges $\mathcal{B}\subseteq\mathcal{E}_{\text{tw}}$, which is $\emptyset$ when generating the initial incumbent, and equal to a branch-and-bound node otherwise. Alg. 2 maintains a stack of multi-agent-labels, or m-labels, for short; in subsequent text, we will always refer to an m-label as an m-label, and never as a label, to avoid confusion with the labels in Section 5.2. An m-label $u=(\gamma,t,\sigma,\vec{b},{n})$ represents a set of partial tours $\{\Gamma^{1},\Gamma^{2},\dots,\Gamma^{{n}}\}$, assigned to agents $1,2,\dots,{n}$, respectively, where $\Gamma^{1},\Gamma^{2},\dots,\Gamma^{{n}-1}$ are tours, and $\Gamma^{n}$ is strictly partial. $u$ represents this set of partial tours as follows:
@@ Appendix E Additional Proofs
Let $\mathcal{F}_{\text{sol}}$ be some $\mathcal{B}$-feasible solution with cardinality at least 1.131313If there is at least 1 target, $|\mathcal{F}_{\text{sol}}|\geq 1$. Even if there are 0 targets, the solution containing a single tour that stays at the depot is $\mathcal{B}$-feasible, and has cardinality 1. Thus, if any $\mathcal{B}$-feasible solutions exist, some $\mathcal{B}$-feasible solution exists with cardinality at least 1. Let $P_{\text{sol}}=(\Gamma^{1},\Gamma^{2},\dots,\Gamma^{|\mathcal{F}_{\text{sol}}|})$ be some permutation of $\mathcal{F}_{\text{sol}}$; note that while $\mathcal{F}_{\text{sol}}$ is a set of tours, $P_{\text{sol}}$ is a sequence of tours. We say that an m-label $(\gamma_{i,j},t,\sigma,\vec{b},n)$ is substitutable into $P_{\text{sol}}$ if the following substitution conditions are satisfied:
@@
The intuition behind the substitution conditions is as follows. By traversing the backpointers of m-label $u=(\gamma_{i,j},t,\sigma,\vec{b},n)$, we will obtain a sequence of partial tours $P_{\text{sol}}^{\prime}=(\Gamma^{{}^{\prime}1},\Gamma^{{}^{\prime}2},\dots,\Gamma^{{}^{\prime}n})$, where $\Gamma^{{}^{\prime}n}$ is strictly partial and the others are tours. Let $l={\textnormal{Len}}(\Gamma^{{}^{\prime}n})$, and let $\Gamma^{n}[l+1:]$ be the suffix of $\Gamma^{n}$ beginning at index $l+1$. Let $\Gamma^{{}^{\prime\prime}n}$ be the tour obtained by concatenating $\Gamma^{{}^{\prime}n}$ with $\Gamma^{n}[l+1:]$, where $\Gamma^{{}^{\prime}n}$ comes first. If the substitution conditions are satisfied, we can construct a $\mathcal{B}$-feasible solution $\mathcal{F}_{\text{sol}}^{\prime\prime}=\{\Gamma^{{}^{\prime}1},\Gamma^{{}^{\prime}2},\dots,\Gamma^{{}^{\prime\prime}n},\Gamma^{n+1},\dots,\Gamma^{|\mathcal{F}_{\text{sol}}|}\}$, where we have essentially substituted partial tours associated with $u$ into $\mathcal{F}_{\text{sol}}$.
We now show that every time the condition on Line 3 is checked, the stack contains an m-label $(\gamma_{i,j},t,\sigma,\vec{b},n)$ substitutable into $P_{\text{sol}}$. Let $\Delta$ be the set of all m-labels that (i) have been popped from the stack, (ii) were not dominated on Line 6, and (iii) are substitutable into $P_{\text{sol}}$. We have two cases:
@@
  • Suppose $n={n_{\textnormal{agt}}}$. Since $u$ is substitutable into $P_{\text{sol}}$, there is a speed-admissible trajectory $\tau_{\textnormal{a}}$} executing $\Gamma^{n}$ and intercepting $\gamma_{i,j}$ at time no earlier than $t$ (substitution condition 3). Let $t^{\#}$ be the time at which $\tau_{\textnormal{a}}$} intercepts $\gamma_{i^{\prime},j^{\prime}}$. The speed-admissibility of $\tau_{\textnormal{a}}$} ensures that ${\textnormal{EFAT}}(\tau_{i}(t),t,\gamma_{i^{\prime},j^{\prime}})\leq t^{\#}$. Furthermore, since each target $i^{\prime\prime}$ not yet intercepted at time $t^{\#}$ by $\tau_{\textnormal{a}}$} is intercepted by $\tau_{\textnormal{a}}$} at some future time, $t^{\#}\leq\max\limits_{j^{\prime\prime}\in\{1,2,\dots,n_{\text{win}}(i^{\prime\prime})\}}{\textnormal{LFDT}}(\gamma_{i^{\prime},j^{\prime}},\gamma_{i^{\prime\prime},j^{\prime\prime}})$. This satisfies condition 5 in the definition of a successor target-window.
@@
Case 2.2.1: Suppose $u^{\prime}$ was dominated by some $u^{\prime\prime}=(\gamma_{i^{\prime},j^{\prime}},t^{\prime\prime},\sigma^{\prime\prime},\vec{b}^{\prime},n)$ in $\mathfrak{D}[\gamma_{i^{\prime},j^{\prime}},\vec{b}^{\prime},n^{\prime})]$ on Line 11. Without loss of generality, let $u^{\prime\prime}$ be an element of $\mathfrak{D}[(\gamma_{i^{\prime},j^{\prime}},\vec{b}^{\prime},n)]$ that is not dominated by any other element of $\mathfrak{D}[(\gamma_{i^{\prime},j^{\prime}},\vec{b}^{\prime},n)]$. $u^{\prime}$ and $u^{\prime\prime}$ both are substitutable into $P_{\text{sol}}$, since all of the following hold:

### v3

@@ 7 Conclusion
In this paper, we introduced BPRC, a new algorithm to find the optimum for the MT-VRP. We demonstrated that it finds the optimum with less computation time than a baseline, particularly when the agents have small capacities. As mentioned in Section 6, pricing is a bottleneck in BPRC, and thus a direction for future work is to employ techniques such as the ng-path relaxation (Baldacci, Mingozzi, and Roberti 2011) to reduce time spent in the pricing problem.
@@ Appendix D Generating Additional Feasible Solutions
This section describes the function used to generate the initial incumbent and generate additional feasible solutions when the RMP is infeasible in Section 5. Given a set of disallowed edges $\mathcal{B}\subseteq\mathcal{E}_{\text{tw}}$, GenerateFeasibleSolution attempts to generate a set of tours solving the MT-VRP while traversing no edge in $\mathcal{B}$. We do so using the DFS in Alg. 2, which attempts to quickly find tours without regard to cost. Alg. 2 takes as input a set of disallowed edges $\mathcal{B}\subseteq\mathcal{E}_{\text{tw}}$, which is $\emptyset$ when generating the initial incumbent, and equal to a branch-and-bound node otherwise. Alg. 2 maintains a stack of multi-agent-labels, or m-labels, for short; in subsequent text, we will always refer to an m-label as an m-label, and never as a label, to avoid confusion with the labels in Section 5.2. An m-label $u=(\gamma,t,\sigma,\vec{b},{n})$ represents a set of partial tours $\{\Gamma^{1},\Gamma^{2},\dots,\Gamma^{{n}}\}$, assigned to agents $1,2,\dots,{n}$, respectively, where $\Gamma^{1},\Gamma^{2},\dots,\Gamma^{{n}-1}$ are tours, and $\Gamma^{n}$ is not a tour. $u$ represents this set of partial tours as follows:
@@ Appendix E Additional Proofs
Let $\mathcal{F}_{\text{sol}}$ be some $\mathcal{B}$-feasible solution with cardinality at least 1.151515If there is at least 1 target, $|\mathcal{F}_{\text{sol}}|\geq 1$. Even if there are 0 targets, the solution containing a single tour that stays at the depot is $\mathcal{B}$-feasible, and has cardinality 1. Thus, if any $\mathcal{B}$-feasible solutions exist, some $\mathcal{B}$-feasible solution exists with cardinality at least 1. Let $P_{\text{sol}}=(\Gamma^{1},\Gamma^{2},\dots,\Gamma^{|\mathcal{F}_{\text{sol}}|})$ be some permutation of $\mathcal{F}_{\text{sol}}$; note that while $\mathcal{F}_{\text{sol}}$ is a set of tours, $P_{\text{sol}}$ is a sequence of tours. We say that an m-label $(\gamma_{i,j},t,\sigma,\vec{b},n)$ is substitutable into $P_{\text{sol}}$ if the following substitution conditions are satisfied:
@@
The intuition behind the substitution conditions is as follows. By traversing the backpointers of m-label $u=(\gamma_{i,j},t,\sigma,\vec{b},n)$, we will obtain a sequence of partial tours $P_{\text{sol}}^{\prime}=(\Gamma^{{}^{\prime}1},\Gamma^{{}^{\prime}2},\dots,\Gamma^{{}^{\prime}n})$, where $\Gamma^{{}^{\prime}n}$ is not a tour and the others are tours. Let $l={\textnormal{Len}}(\Gamma^{{}^{\prime}n})$, and let $\Gamma^{n}[l+1:]$ be the suffix of $\Gamma^{n}$ beginning at index $l+1$. Let $\Gamma^{{}^{\prime\prime}n}$ be the tour obtained by concatenating $\Gamma^{{}^{\prime}n}$ with $\Gamma^{n}[l+1:]$, where $\Gamma^{{}^{\prime}n}$ comes first. If the substitution conditions are satisfied, we can construct a $\mathcal{B}$-feasible solution $\mathcal{F}_{\text{sol}}^{\prime\prime}=\{\Gamma^{{}^{\prime}1},\Gamma^{{}^{\prime}2},\dots,\Gamma^{{}^{\prime\prime}n},\Gamma^{n+1},\dots,\Gamma^{|\mathcal{F}_{\text{sol}}|}\}$, where we have essentially substituted partial tours associated with $u$ into $\mathcal{F}_{\text{sol}}$.
We now show that every time the condition on Line 3 is checked, the stack contains an m-label $(\gamma_{i,j},t,\sigma,\vec{b},n)$ substitutable into $P_{\text{sol}}$. Let $\Delta$ be the set of all m-labels that (i) have been popped from the stack, (ii) were not dominated on Line 5, and (iii) are substitutable into $P_{\text{sol}}$. We have two cases:
@@
  • Suppose $n={n_{\textnormal{agt}}}$. Since $u$ is substitutable into $P_{\text{sol}}$, there is a speed-admissible trajectory $\tau_{\textnormal{a}}$} executing $\Gamma^{n}$ and intercepting $\gamma_{i,j}$ at time no earlier than $t$ (substitution condition 3). Let $t^{\#}$ be the time at which $\tau_{\textnormal{a}}$} intercepts $\gamma_{i^{\prime},j^{\prime}}$. The speed-admissibility of $\tau_{\textnormal{a}}$} ensures that ${\textnormal{EFAT}}(\tau_{i}(t),t,\gamma_{i^{\prime},j^{\prime}})\leq t^{\#}$. Furthermore, since each target $i^{\prime\prime}$ not yet intercepted at time $t^{\#}$ by $\tau_{\textnormal{a}}$} is intercepted by $\tau_{\textnormal{a}}$} at some future time, $t^{\#}\leq\max\limits_{j^{\prime\prime}\in\{1,2,\dots,n_{\text{win}}(i^{\prime\prime})\}}{\textnormal{LFDT}}(\gamma_{i^{\prime},j^{\prime}},\gamma_{i^{\prime\prime},j^{\prime\prime}})$ for all such targets $i^{\prime\prime}$. This satisfies condition 5 in the definition of a successor target-window.
@@
Case 2.2.1: Suppose $u^{\prime}$ was dominated by some $u^{\prime\prime}=(\gamma_{i^{\prime},j^{\prime}},t^{\prime\prime},\sigma^{\prime\prime},\vec{b}^{\prime},n)$ in $\mathfrak{D}[\gamma_{i^{\prime},j^{\prime}},\vec{b}^{\prime},n)]$ on Line 11. Without loss of generality, let $u^{\prime\prime}$ be an element of $\mathfrak{D}[(\gamma_{i^{\prime},j^{\prime}},\vec{b}^{\prime},n)]$ that is not dominated by any other element of $\mathfrak{D}[(\gamma_{i^{\prime},j^{\prime}},\vec{b}^{\prime},n)]$. $u^{\prime}$ and $u^{\prime\prime}$ both are substitutable into $P_{\text{sol}}$, since all of the following hold: